OpenStudy (anonymous):

The measures of 2 complementary angles are 16z-9 and 4z+3. find the measures of these angles.

OpenStudy (anonymous):

Add them together and make them equal to 90 degrees

OpenStudy (anonymous):

16z-9+4z+3=90 16z+4z=96 20z+96 z=4.8

OpenStudy (aroub):

16z-9+4z+3=90 20z-6=90 20z=96 z=96/20 z=24/5 z=4.8

OpenStudy (anonymous):

Complementary means they add up to 90 degrees Supplementary means they add up to 180 degrees A good way to remember that is "C" comes before "S" in the alphabet and 90 comes before 180.

OpenStudy (anonymous):

Z=4.8